Dry ice is frozen carbon dioxide. If you have 1.8
Answer #1

Dry ice is frozen CO₂
The molecular mass of CO₂ is 12.0 + 2·16.0 = 44.0 g/mol

Ideal Gas Law: PV = nRT --------------(1)
R = 8.314 J/K·mol

1.8 kg = 1800 g of CO₂ is
n = (1800/44.0) mol = 40.91

0ºC = 273.15 K; T =20ºC = 293.15 K

P = 1 atm = 1.013·10⁵ N/m²

From (1),

V = nRT/P = (40.91*8.314*293.15)/( 1.013·10⁵) = 0.98m^3